### v1

# Let’s Influence Algorithms Together: How Millions of Fans Build Collective Understanding of Algorithms and Organize Coordinated Algorithmic Actions

Abstract.
Previous research pays attention to how users strategically understand and consciously interact with algorithms but mainly focuses on an individual level, making it difficult to explore how users within communities could develop a collective understanding of algorithms and organize collective algorithmic actions. Through a two-year ethnography of online fan activities, this study investigates 43 core fans who always organize large-scale fans’ collective actions and their corresponding general fan groups. This study aims to reveal how these core fans mobilize millions of general fans through collective algorithmic actions. These core fans reported the rhetorical strategies used to persuade general fans, the steps taken to build a collective understanding of algorithms, and the collaborative processes that adapt collective actions across platforms and cultures. Our findings highlight the key factors that enable computer-supported collective algorithmic actions and extend collective action research into large-scale domain targeting algorithms.
1. Introduction
In December 2020, Weibo, one of China’s largest social media platforms, created a separate trending section for entertainment after fan communities repeatedly manipulated the platform’s trending algorithm to promote celebrities and entertainment topics. This change aimed to prevent entertainment content from overshadowing more relevant public news and significant events on the main trending list (Entertainment, 2022; Zhang et al., 2023; Net, 2023). Similarly, in 2023, South Korea’s leading music streaming platform, Melon, modified its ranking algorithm to exclude muted plays, a tactic K-pop fans used to artificially boost their idols’ rankings. In 2022, this manipulation generated an estimated $33.2 million in revenue (Dredge, 2023). We define this phenomenon as ”fan-driven collective algorithmic action,” wherein large-scale fans organize and mobilize to strategically alter algorithmic outputs on social media, aiming to enhance the visibility of their favorite idols. This collective effort encompasses a range of tactics that fans believe can influence algorithms, such as promoting specific topics to trend or engaging in repetitive muted video plays to elevate their idols’ rankings (Zhang et al., 2023; Xiao and Zheng, 2023).
Past HCI research has explored how various kinds of users strategically develop an understanding of platform algorithms and subsequently negotiate with these algorithms (DeVito et al., 2017; Cotter, 2019; Karizat et al., 2021; DeVos et al., 2022). For instance, Cotter discussed how Instagram influencers develop an instrumental interaction of algorithms to increase their visibility by discerning the operational principles of these algorithms (Cotter, 2019). Yu et al. discovered that delivery workers formed micro-scale algorithmic communities by establishing private social media groups for mutual support. In these groups, they collaboratively developed a folk understanding of algorithms and shared insights into algorithmic functioning and resistance strategies (Yu et al., 2022). These studies have investigated how individuals or small groups of individuals make use of or resist algorithms and progressively develop their understanding of algorithms. However, at the collective level, particularly when considering larger groups of millions or even tens of millions, it remains unclear to HCI scholars how communities collectively understand algorithms and how collective actions towards algorithms could be developed and organized.
In contrast to the academic community’s lack of research on large-scale collective action targeting algorithms, such behaviors are on the rise and are having a significant impact on both platforms and society (Bonini and Treré, 2024; Treré and Bonini, 2024). Citizens are increasingly dissatisfied with their passive role in the face of algorithms and are seeking to enhance their influence through algorithmic activism (Velkova and Kaun, 2021; Bonini and Treré, 2024). For instance, political groups have previously engaged in Google bombing by utilizing predictions of Google’s algorithms to develop strategies for designing websites, thereby influencing Google search outcomes (Gillespie, 2019). In 2018, activists manipulated Google’s algorithm and organized a large group of political allies on Reddit, massively upvoting Trump-related posts, which caused Google’s ranking algorithm to associate the term ‘idiot’ with his image (Haynes, 2018; Gillespie, 2019).
As previous algorithm scholars have pointed out, a significant trend in digital platforms is that more and more groups are using algorithms as tools for social movements and for gaining visibility (Bonini and Treré, 2024; Treré and Bonini, 2024; Gillespie, 2019). These groups amplify their influence over algorithms and hijack platform algorithms to achieve their own goals (Treré and Bonini, 2024). However, there is a lack of research in academia on how these groups mobilize and organize large-scale actions targeting algorithms.
Our research uses large-scale (millions) fan-driven, computer-supported collective actions targeting algorithms (occurring across multiple countries and platforms) as a case study to expand the understanding of HCI and related scholars regarding such large-scale collective algorithmic actions.
Specifically, in this paper, we aim to address the following research questions:
RQ1: What motivates fans to mobilize collective algorithmic actions?
RQ2: How do fans build a collective folk understanding of algorithms?
RQ3: How do fans organize and execute collective algorithmic actions?
To address these questions, we conducted a two-year ethnographic study (Harrison, 2018) of fan-driven collective algorithmic actions from 2021 to 2023. Previous research on fan communities has shown that collective algorithmic actions are often led by a small number of core fans. This demonstrates a hierarchical structure within fan communities, where core fans have the authority to mobilize large numbers of general fans for collective actions. In contrast, general fans typically occupy a more peripheral position regarding power within these communities (Zhang et al., 2023; Zheng and Xiao, 2023; Wu, 2021).
Therefore, we observe 43 core fans and their corresponding general fan groups in collective fan algorithmic actions in a two-year ethnographic study. Additionally, we conducted one-on-one interviews and focus groups with the 43 core fans. On top of that, we interviewed 15 general fans to supplement their perspectives on collective algorithmic action. All the interviewed core fans have extensive experience organizing collective algorithmic actions and hold central roles in their communities, managing fan clubs with over 100,000 followers and maintaining personal social media accounts with more than 5,000 followers each. The general fans, with at least two years of participation, follow the directives of core fans in these actions but lack decision-making authority and influence over strategy, focusing instead on execution and support.
Throughout this longitudinal observation, we studied how fan communities, as large-scale collectives, develop, refine, and test their understanding of algorithms, converting this understanding into collective algorithmic actions. We also investigate how a small number of core fans persuade, guide, and coordinate millions of general fans to participate in these daily collective efforts to influence algorithms.
Our work’s key contributions to the HCI community are as follows:
Through a two-year ethnographic study, complemented by semi-structured interviews and focus groups, we conducted an in-depth analysis of how millions of fans orchestrate influential collective algorithmic actions.
We advanced folk theories of algorithms from an individual dimension to a collective understanding and explored how these folk theories evolve into well-defined, context-specific algorithmic action. We also extend computer-supported collective action studies to the field of algorithms.
We explored the potential for large-scale, computer-supported collective action by examining fan-driven algorithmic efforts capable of mobilizing millions of users. Our study, encompassing multiple fan communities, nations, and platforms, demonstrates not only the broad applicability of our theory to future large-scale collective actions but also potential strategies for overcoming large-scale collective action dilemmas.
2. Related Work
2.1. Folk Theories of Algorithm
Folk theories of algorithms are user-generated explanations that help people understand how algorithms work and influence them (Ytre-Arne and Moe, 2021; Eslami et al., 2019). Users develop various folk understandings of how algorithmic systems operate and produce different outcomes, considering that these systems remain opaque to users (Karizat et al., 2021; Shen et al., 2021; Li et al., 2023). Research has explored how users develop these theories through their direct interactions with algorithms (Cotter, 2019; Eslami et al., 2019), algorithmic gossip (Bishop, 2019; Ticona and TSAPATSARIS, 2023; Zhang et al., 2023), and social interactions within their communities (Mayworm et al., 2024). Through these means, users come to understand that social media information and interactions are driven by algorithmic systems and even learn some details about how these algorithms work. However, they still do not know the specific mechanisms behind the algorithms. This lack of full understanding leads them to develop their own folk theories for understanding their personal experiences with algorithms (Shen et al., 2021; Eslami et al., 2019; Li et al., 2023).
HCI studies have examined how users perceive various aspects of algorithms, such as how they affect user visibility on social media (Alvarado and Waern, 2018; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022), the processes behind content moderation (Mayworm et al., 2024; Moran et al., 2022), how algorithms facilitate user matching on dating apps and how news is curated by algorithms (Ngo and Krämer, 2022) These previous studies reveal how users build and refine their understanding of algorithms based on their day-to-day experiences and observations. In this process of testing and updating, users can develop algorithmic folk theories that are sufficiently intricate and flexible (Karizat et al., 2021). As they interact with algorithms and refine their understanding through ongoing experiences and observations, these folk theories become more nuanced. Users adjust their beliefs based on new information and changing interactions, allowing their folk theories to evolve through an iterative process in response to varying algorithmic behaviors and outcomes (DeVito, 2021; Zhang et al., 2023).
Despite this, these folk theories of algorithms are often not entirely accurate because algorithms remain a ”black box” to users, who typically lack detailed knowledge of the platform’s algorithms or system design. However, these studies suggest that users have the potential to actively drive the testing and refinement of algorithms, gradually improving the reliability of their folk theories. This user-driven approach can enhance algorithmic transparency and understandability from the user’s perspective (Shen et al., 2021; Eslami et al., 2019).
2.2. Computer-Supported Collective Action
Olson defines collective action as efforts by a group to achieve common goals, such as lobbying, enhancing competitiveness, or engaging in protests (Olson Jr, 1971). This action is driven by collective interests, where benefits are shared and non-excludable, leading to the free-rider problem, where individuals benefit without contributing. Olson argues this issue is more severe in larger groups due to difficulties in organization, higher persuasion costs, and reduced individual incentives (Olson Jr, 1971). Consequently, Olson (Olson Jr, 1971) notes that collective action often results in ”collective inaction,” where groups fail to mobilize despite shared interests.
Effective collective action requires an organizational mechanism to coordinate and incentivize participants, such as providing incentives, establishing structures, or guiding through leadership (Olson Jr, 1971). Factors like group size, proximity, meeting frequency, uniformity, and organizing technologies also play crucial roles (Olson Jr, 1971; Elster, 1985). Elster (Elster, 1985) further highlights collective consciousness as key to addressing the free-rider problem. This awareness connects participation to personal and collective identity, often fostered by experienced individuals and opinion leaders.
Recent HCI researchers have increasingly focused on computer-supported collective action due to its potential to generate significant attention on social media and create real-world impacts (Shaw et al., 2014; Salehi et al., 2015; Bennett and Segerberg, 2013; Matias, 2016; Wu et al., 2022; Huang et al., 2015; Yasseri et al., 2016). Salehi et al. (Salehi et al., 2015) found that the web has the potential to enable distributed collectives to address shared issues. By examining the collective actions of over 100 Amazon Mechanical Turk workers, they explored the dynamics of computer-supported collective efforts. They introduced the concept of tactical publics, highlighting the transition from public deliberation to tactical actions and identifying stalling and friction as significant challenges in this process. Matias (Matias, 2016) examined how 2,278 online community leaders on Reddit coordinated collective actions to influence platform operators. Their efforts led to millions of readers being restricted from accessing significant portions of Reddit and persuaded the company to negotiate to address their demands.
Shaw et al. (Shaw et al., 2014) proposed five stages in computer-supported collective action:
Stage 1. Identify a problem and others who care about it. Recognize key issues affecting a specific public and find others who share the concern.
Stage 2. Generate, debate, and select viable solutions. Focus on creating and choosing practical solutions once aligned with a committed group.
Stage 3. Coordinate and prepare for action. Plan the necessary steps, assign leadership, define tasks, set milestones, and clarify timelines and responsibilities.
Stage 4. Take action. Maintain coordination to adapt to new information and changing circumstances.
Stage 5. Assess, document, and follow up. Evaluate, document, and review outcomes to prepare for future work efficiently.
However, it’s important to note that many collective actions on social media fail to achieve their goals; for instance, over 99% of Change.org petitions aimed at improving privacy protections never reach ”victory” status (Wu et al., 2022; Huang et al., 2015). To conclude, many of these efforts fail because they bypass key stages in the model proposed by Shaw et al. (Shaw et al., 2014), such as effective issue framing, consensus-building, and the structured coordination of collective actions. Users also often lack the specific computational tools to collectively action.
Previous HCI research has rarely addressed successful computer-supported collective actions. Even when it does, it seldom involves actions on the scale of millions of people (Salehi et al., 2015; Matias, 2016), particularly those specifically targeting algorithms. Our work uses large-scale, cross-platform, and cross-cultural fan algorithmic collective action as a case study to evaluate the iterative process of such computer-supported collective action development. We discuss how collectives develop a unified understanding of algorithms, translate this understanding into action, and effectively coordinate efforts across multiple platforms and cultures to address the organizational challenges faced in previous collective action. Additionally, we explore the resources and theoretical frameworks required for organizing computer-supported collective actions, focusing on those targeting algorithms.
2.3. Transformative Fandom
Previous HCI research has investigated fan communities’ transformative potential, focusing on their communal and collective aspects. Wang et al.(Wang et al., 2024) explored how privacy theory extends from individuals to collectives by examining fan self-disclosure practices in online communities. They found that fans engage in value construction, trust-building, and personal disclosure, fostering a supportive collective environment. Park et al.(Park et al., 2021) studied how online fan communities effectively collaborate on causes beyond their original purpose. They revealed that fans use shared values, teamwork, and established infrastructure to achieve large-scale collaboration and build mutual trust. For instance, ARMY, BTS’s fanbase, organized the #MatchAMillion campaign in 2020, raising over one million USD for Black Lives Matter. This transformative aspect has made fans a key area of study in transformative community and self-organizational behavior (Zhang et al., 2023; Wang et al., 2024; Jenkins, 2012; Dym and Fiesler, 2018; Fiesler et al., 2016).
Our case study focuses on Korean Popular Music (K-pop) and Chinese Popular Music (C-pop) idol fans, whose communities are closely linked with the idol industry (Jenkins, 2012). Fans derive emotional satisfaction from idols (Yin, 2020) and contribute to their success through online promotion and gifting (Yin, 2020; Stanfill, 2013; Yang, 2009). These fan communities are well-structured (Wu, 2021). Wu (Wu, 2021) described how fan clubs, managing various fan activities, are central to this structure. In Wu’s framework (Wu, 2021), fan clubs and their departments are core fans, while general fans support idols based on their guidance. Zheng and Xiao (Zheng and Xiao, 2023) emphasized the core fans’ authority, noting that general fans support them through actions like retweeting and positive commenting and engage in significant data labor under their direction.
Recent research focuses on fans strategically engaging with platform algorithms and developing collective algorithmic actions to enhance their influence and idols’ visibility (Zhang et al., 2023; He and Li, 2023; Zheng and Xiao, 2023; Zhang and Negus, 2020). Yin (Yin, 2020) found that fans willingly engage in unpaid data labor intended to influence algorithms and increase idol visibility. This includes basic activities like liking, commenting, and reposting, as well as promoting hashtags and countering negative comments (Yin, 2020; Zheng and Xiao, 2023; Zhang and Negus, 2020). For example, Chinese fans have normalized data labor on platforms like Weibo, seeking to optimize their strategies based on algorithmic rules. Yin (Yin, 2020) observed that collective emotions within fan communities strongly motivate this labor. Long-term data labor has shifted the meaning of fan practices on social media from consumption and interaction to focusing on data contribution for boosting idol visibility. Zhang et al. (Zhang et al., 2023) examined how fans perceive the link between their algorithm manipulation activities and idol visibility, using platform reactions as a basis. They explored how fans integrate computational tools with manual labor and found that algorithmic imaginaries spread from core groups to the wider fan base through hierarchical and vertical processes.
Although these studies have explored the transformative potential of fans and the phenomenon of fans developing collective actions targeting algorithms, research is still lacking on how these large-scale collective actions are organized within fan communities. While earlier studies have provided a theoretical foundation for collective algorithmic imaginaries among fans, our work primarily contributes by examining how these collective algorithmic understandings undergo iterative development through collective efforts, extend to actionable levels, and become integrated into everyday practices across the community.
3. Method
In this study on fans’ collective algorithmic actions, we define ”core fans” as those who hold organizational power within the fan community, possess significant reputations, and directly influence the organization of collective fan actions. In contrast, ”general fans” are those who participate in collective algorithmic actions but do not directly engage in core decision-making or shape the overall strategy of the actions; their primary role is to execute tasks as directed.
More specifically, we conducted an in-depth two-year ethnographic study from 2021 to 2023, focusing on interactions with 43 core fans in China. The study not only involved observing these 43 Chinese core fans but also included tracking and analyzing the daily activities of their respective fan clubs and the broader general fan groups, all with their consent. To supplement the empirical data, we also conducted one-on-one interviews with 15 general fans from their fan clubs.
The study received approval from the university’s ethics review board. Throughout the research process, the researcher maintained transparency regarding their role and obtained informed consent from all participants. We consistently prioritized listening to the voices of participants, critically reflecting on our positionality as researchers, and conscientiously examining potential biases.
Throughout these two years, we engaged with these Chinese core fans biweekly to gather their reflections on their daily collective algorithmic actions. These reflections encompassed both the internal strategies within the core fan group, as well as their influence on the general fan group. This ethnographic method allowed us to understand the complex behavioral patterns of fans in collective algorithmic actions and how they coordinate across different levels of fan groups to drive successful collective action.
3.1. Participating Role
Inspired by previous HCI and fan studies (Thach et al., 2023; Zheng and Xiao, 2023; Joshi et al., 2024), this research adopts an ethnographic method with a participant-observer approach. We reached out to 60 core fans, and ultimately, 43 of them agreed to participate in the study on the condition of complete anonymity. With their consent, two of our research team members immersed themselves in the fans’ daily algorithm-related activities, which included internal discussions among core fans about algorithm action, designing a rhetorical language for persuading general fans, deciding which social media activities required collective algorithmic action, evaluating the effectiveness of these actions, and iterating on the processes. These researchers actively participated in these core fans’ internal discussions.
Furthermore, these two researchers also engaged in the algorithmic practices typical of general fans, dedicating 1 to 2 hours daily, from Monday to Friday, to perform routine activities such as repetitive liking, reposting, and commenting, and adjusting specific strategies across different platforms. The platforms where our researchers conducted these algorithmic practices included YouTube, Weibo, X (formerly Twitter), Douyin, TikTok, Xiaohongshu, and other popular platforms where fan activities are active. Additionally, they actively joined and observed the online communities of general fans within the fan clubs corresponding to these 43 core fans, participating in their discussions.
3.2. Participants: Core Fans (N=43), General Fans (N=15)
The 43 Chinese core fans (CF1 to CF43) we interviewed all actively participated in organizing daily fan algorithmic actions. In addition, each organized over 30 specialized collective algorithmic actions targeting specific events, such as brand campaigns involving their idol, showcasing their extensive experience in coordinating such activities. The platforms impacted by the algorithmic actions discussed in these interviews included major fan activity platforms like YouTube, Weibo, X (formerly Twitter), Bilibili, Douyin, TikTok, Xiaohongshu, WeChat Channel and various music charts. Among the Chinese core fans we interviewed, 4 (CF1 to CF4) managed fan clubs with 100,000 to 200,000 followers on their social media accounts, 21 (CF5 to CF25) managed clubs with 200,000 to 500,000 followers, 11 (CF26 to CF36) managed clubs with 500,000 to 1,000,000 followers, and 7 (CF37 to CF43) managed clubs with more than 1,000,000 followers. According to our participants, the follower count on a fan club’s social media account is a more accurate indicator of the number of people involved in collective algorithmic actions compared to the follower count on the idol’s own social media account. In their view, following the idol might indicate an appreciation for the idol personally, but it doesn’t necessarily imply active participation in algorithmic actions. In contrast, those who follow a fan club are much more likely to be highly motivated to engage in fan-driven algorithmic activities. Therefore, the data from these fan clubs can assist HCI researchers in understanding the scale of their organized algorithmic actions. These interviewed Chinese core fans could directly participate in and influence their fan clubs’ decisions and judgments regarding algorithmic folk understanding and collective algorithmic actions. Additionally, all the core fans interviewed have their own social media account besides their fan club, while each had over 5,000 followers on social media platforms, with a monthly average readership exceeding 2 million, and their social media content primarily focused on their idol and fan community, thus affirming their core fans status (sufficient influence driven by their own fan identity).
Fifteen general fans (GF1 to GF15) were voluntarily recruited from the fan clubs of the previous 43 core fans. Each has over 5 years of fan experience and more than 2 years of participation in collective data actions. During the two years of our ethnographic study, they were actively engaged in algorithmic practices almost every day. They are distant from the decision-making groups of core fans and do not participate in the writing of algorithm tutorials or related discussions. Instead, they engage in the everyday algorithmic actions typical of general fans.
3.3. Data Collection
3.3.1. Online and on-site observation
We joined core fan groups on various social media platforms, such as WeChat, Weibo, and WhatsApp, where algorithms for organizing and coordinating actions are discussed, to understand how core fans perceive algorithms and how they extend this understanding to collective actions. Additionally, the two lead researchers participated in 22 offline gatherings of core fans to better grasp their perspectives and actions. We also joined the social media groups of general fans, focusing on discussions surrounding folk understanding of algorithms and the coordination of collective algorithmic actions. In these instances, the researchers were not passive participants; rather, they engaged in participatory observation as part of the fan community, fully immersing themselves in these algorithm-related fan activities (Zheng and Xiao, 2023). This participatory observation approach follows an insider’s perspective, allowing the two lead authors to document and discuss the activities of core fans and the virtual or physical spaces they visited, thereby enabling interactive verification of data collection with other interviewees (Zheng and Xiao, 2023). These observations spanned from 2021 to 2023.
3.3.2. Interviews with core fans and general fans
Semi-structured interviews often effectively complement ethnographic observation data (Thach et al., 2023). Therefore, we conducted semi-structured interviews with the 43 core fans we observed and an additional 15 general fans from their fan clubs. These interviews were conducted online following the ethnographic observations to gather their views on collective algorithmic actions. All interviews were recorded and transcribed by researchers. The semi-structured interviews with core fans focused on their motivations for organizing collective algorithmic actions, challenges, the distinctions between different algorithmic actions, the iteration processes, and the reasoning behind these iterations. These interviews for core fans ranged from 1 to 3 hours, depending on the core fans’ availability. For the general fans, the semi-structured interviews explored their motivations for participating in collective algorithmic actions, their perceptions of the effectiveness of these actions, their views on the persuasive rhetoric and algorithmic action tutorials organized by core fans, and their assessment of the overall performance of the fan communities and fan clubs. These interviews for general fans ranged from 1 to 1.5 hours.
3.3.3. Focus group with core fans
After conducting observations and interviews, we held three focus group sessions, each with an average of five participants lasting approximately two and a half hours, to gather a collective perspective on algorithmic understanding and actions among core fans. In HCI research, focus groups are a common complement to ethnography, enabling the exploration of group dynamics, collective reasoning, and the communal construction of knowledge. Unlike individual interviews, focus groups facilitate interactions among participants, bringing to light shared concerns, common strategies, and differing viewpoints—all crucial for understanding the collective behaviors that drive fan-led algorithmic practices.
3.4. Data Analysis
Through a thematic analysis approach (Braun and Clarke, 2012, 2019), two researchers conducted a comprehensive qualitative analysis of the field notes, interview transcripts, and focus group discussions. By applying open coding(Braun and Clarke, 2012, 2019), the researchers could break down the data into discrete elements, making it possible to identify themes and relationships that might not have been immediately apparent. This collaborative process of discussion, reflection, and refinement also helped to enhance the reliability of the analysis. We did not calculate inter-coder reliability between the two researchers. Instead, we adhered to the principles of reflexive thematic analysis(Braun and Clarke, 2012, 2019), engaging in ongoing critical reflection and open discussions to address any consensus or discrepancies that arose during the coding process. After completing the coding, the research team conceptualizes higher-level themes from these codes, employing the affinity diagram method to organize the emerging themes. This process ultimately resulted in the identification of first-level themes, second-level themes, and third-level themes. Our findings are detailed in Section 4, with the section headings aligning closely with the second and third-level themes. A comprehensive list of all second and third-level themes can be found in the appendix ( Table 1).
3.5. Ethics
We required all participants to sign informed consent forms and consistently disclose our researcher identity throughout the study. Participants were given the option to withdraw from the research at any time. During the ethnographic observations, we adhered to ethical requirements by not directly collecting materials; the researchers compiled secondary field notes based on their observations for research purposes. In the interviews and focus group sessions, we sought participants’ consent and encouraged them to ask any questions they had about the research. Throughout data analysis and manuscript preparation, we implemented strict anonymization procedures to protect participants’ identities.
4. Results
Algorithmic labor is a prevalent phenomenon among fan communities, primarily involving fans engaging in intentional actions through platform algorithms to enhance the visibility of their idols. This includes increasing repost numbers on social media platforms, video platform views, and music platform rankings. These actions are often repetitive, named “algorithmic labor” in fandom (Zheng and Xiao, 2023; Zhang et al., 2023). In Chinese fandom, fans sometimes refer to this behavior as ”data labor” because they are directly engaged with various forms of data, such as shares and views, while the algorithm remains invisible.
4.1. “All for Our Idol”: Recognizing the Necessity of Collective Algorithmic Action (RQ1)
The first step in fan-driven collective algorithmic action is recognizing the necessity of such action. Core fans and general fans, when asked why fans engage in algorithmic action, unanimously see it as a key way to support their idols. They clearly understand that the behavior of fans on social media has a significant impact on the visibility of their idols, especially when influenced by recommendation algorithms. They have made a conscious connection between how fans engage with social media and the level of exposure their idols receive. By closely examining the various types of data generated by fan activities, they have been able to distinguish how different forms of social media engagement—such as likes, shares, comments, and other interactions—affect the visibility and prominence of their idols within the algorithmically driven visibility of social media platforms.
In the view of the core fans we interviewed, recognizing the impact of fan-platform interactions on their idols, as well as understanding the specific roles of different forms of interaction, is the first step in engaging in algorithmic action. This awareness makes them realize that taking algorithmic action is necessary and feasible. CF43 explains to us: ”First, we need to understand that developing algorithms is beneficial for idols, right? And this benefit is quite significant; otherwise, we wouldn’t be working so hard on this. Second, we must be clear about the practical impact of our approach; we can’t just run around like headless chickens.” Next, we will explore how these fans perceive and interpret various fan behaviors on social media.
4.1.1. The Importance of Data Performance in Idols’ Commercial Collaborations
Before organizing collective algorithmic actions, people must first recognize that such actions are essential and necessary. Fans believe that an idol’s data performance on social media significantly influences their commercial opportunities. CF38, a fan club leader, explains that brands don’t just subjectively choose celebrities as ambassadors; objective data on social media is crucial. Brands work with data analysis firms and social media or e-commerce platforms for relevant data. This data significantly influences their choice of an idol.
CF14, a fan club administrator, said that an idol partnering with a top brand gains not only significant monetary benefits but also promotional support from the brand. This includes invitations to events and shows sponsored by the brand.
Certainly, impressive ”data performance” is not limited to securing brand endorsements. CF42, a devoted fan with 15 years of fan experience, illustrated this point with one of her 2021 posts on Weibo, which stated that not only commercial brands but also reality shows and galas consider ”data performance” when selecting their guests. While these platforms might not always seek assistance from professional data companies, given their less commercial nature, they still rely on intuitive and direct data metrics as significant factors in choosing celebrities as guests. Finally, this post garnered over 5,000 likes, 600 reposts, and 500 comments on Weibo, earning widespread recognition from the fan community.
All core fans and interviewees in this study agree that ”data performance” is vital for an idol’s commercial appeal. CF32 explains, “Brands value an idol’s data for gauging public visibility, reputation, and fan loyalty. Better data suggests more advantages for the idol.” She highlights fans’ role, noting, “To improve our idol’s chances, we focus on improving their data metrics.”
4.1.2. Perceived Relationship Between Fan-Driven Collective Algorithmic Action and Social Media Algorithms
After recognizing that such action is necessary, people must also understand that their efforts, especially large-scale collective actions, can indeed influence algorithms in their perception. In the views of fans, their actions on social media are likely to influence, or even manipulate, the algorithms of these platforms.
Fans first distinguish between data that can be achieved without manipulating the algorithm and data performance that requires algorithmic intervention, thereby establishing the necessity for fans to take action against the algorithm. CF29, a core fan responsible for organizing fan-driven data performances, outlined the three primary dimensions of core data: (1) monetary-related data, (2) data pertaining to votes and rankings, and (3) social media engagement data.
The monetary data involves sales volumes linked to the idol, encompassing album sales, magazine sales, and endorsed product sales. This data typically reflects the commercial influence of the idol’s core fanbase, indicating their purchasing power and the potential for a brand to recoup costs from idol endorsements. The data pertaining to votes and rankings gauge the willingness of the idol’s core fanbase to devote time and effort, potentially acting as a ’water army’—a term for those hired to manipulate online content—for the brand without compensation.
CF29 noted, ”When our idol is chosen as a brand ambassador, our data performance efforts indirectly benefit the brand, as we engage in activities like reposting, liking, and commenting on specific posts or videos.” She added, ”An idol’s social media engagement data is closely tied to their public visibility and reputation. Typically, better engagement data equates to stronger publicity.” CF33, an administrator of a data team in fandom, further explained that in the realm of reality shows, an idol’s public exposure is often seen as indicative of potential program ratings, making publicity data a critical factor in the invitation process for reality shows.
The first two types of data, concerning monetary contributions and votes or rankings, gauge the engagement of an idol’s core fans. Fans can elevate their idol’s data through straightforward methods like repeated voting and financial expenditure. However, the third type of data, social media engagement data, cannot be manipulated directly. This is because social media platforms employ complex algorithms to filter out artificial publicity data and focus on acquiring more genuine engagements. Therefore, to enhance this type of data performance, fans must carefully devise strategies to influence the algorithm.
CF29 provided an example regarding the order arrangement within comments on Weibo. According to CF29, Weibo utilizes a metric termed ”heat” for its specific algorithm to rank comments. While ”heat” is not directly visible to users, it plays a crucial role in how the platform orders comments. The exact components of ”heat” are not publicly known.
Based on fans’ observations, they believe it encompasses the number of likes, the volume and likes of nested comments, the commenter’s account quality, and user interaction. Fans liken the dense layers of comments, which can number in the hundreds or thousands, to a ’building,’ with nested comments (comments on comments) being a ’building within a building.’ However, Weibo does not disclose the precise criteria for calculating ”heat.”
This lack of transparency arises because fans often strategically like or comment on posts favorable to their idols to boost those comments’ ”heat,” thereby increasing their visibility. This tactic, known as ”controlling comments under posts,” which means conscious control of the comments section by fans, is seen by the public as significantly impacting their online experience, leading to calls for Weibo to mitigate the influence of fan-driven ”controlling comments.” Consequently, Weibo has revised its rules for ”heat” multiple times. Despite these changes, fans continually adapt, tweaking their strategies to align with the new algorithmic logic. This dynamic results in a delicate balance between the platform’s efforts to maintain authentic engagement and the fans’ endeavors to promote their idols.
Fans recognize that their various actions across multiple platforms can significantly impact their idol’s digital visibility, prompting them to continuously expand the scope of their actions. In addition to ”controlling comments” beneath Weibo posts, several other digital arenas have emerged as critical battlegrounds in the interaction between fans and algorithms. These include trending topics on social media, popularity rankings on video platforms, view counts, and algorithmic recommendations. Each area represents a significant opportunity for fans to influence the online visibility and perceived popularity of their favorite celebrities. By engaging with these platforms, fans attempt to navigate and sometimes manipulate algorithmic systems to boost their idol’s digital presence and success.
4.2. Fan-Driven Folk Theories: The Foundation for Guiding Collective Algorithmic Actions (RQ2)
The second step in collective algorithmic action requires a clear understanding of the different mechanisms by which algorithms operate on various platforms and even across different countries. Although this knowledge remains at a folk level, it must be detailed enough to effectively guide their actions. Next, we will demonstrate how core and general fans collaborate to develop a comprehensive understanding of platform algorithms.
CF26, a core fan within the Chinese fandom of a K-pop idol, emphasized that their primary role in collective algorithmic action is to mobilize general fans to engage in data labor. According to CF26, this task requires core fans to ”decode social media algorithms” and guide general fans in manipulating specific data metrics, such as shares and comments, to improve the idol’s overall data performance. Core fans must first understand how the algorithm functions and how to optimize content for recommendations. Only then can they make the process of algorithm manipulation understandable and executable, enabling even fans without technical expertise to participate effectively and follow the guidance provided.
4.2.1. Fan-Driven Methods in Understanding Algorithms.
Before launching collective actions targeting algorithms, people must have a method to thoroughly understand the algorithms, allowing them to perceive whether their actions can effectively influence them and adjust their strategies accordingly based on this assessment. Core fans use the term ”decode” to describe the initial step in developing their understanding of algorithms. They see the algorithm as a black box; users are unaware of its exact operational logic and can only decode it through continuous trial and error. Eventually, this user-driven process creates a set of tutorials that can genuinely influence the algorithm’s output in fans’ perception.
CF26 explained that the process of algorithm decoding by core fans is largely based on folk empirical methods, as they lack direct support from the platform and cannot analyze it from a code or computational perspective. Their approach typically involves trial and error to develop the most effective strategies.
To develop a clear understanding of the algorithm, initially, core fans often gather basic tutorials from other fan groups, as these groups also engage in decoding algorithms and instructing their fans. This is because all fan groups are actively decoding various social media algorithms to increase the visibility of their idols. To engage more general fans, these algorithm manipulation tutorials are always made public. As previously mentioned, there is already a consensus among core and general fans that various interactions on social media significantly impact their idols. Then, it is important to understand which specific interactions can most effectively leverage recommendation algorithms to maximize the visibility of their idols.
By observing patterns in algorithm behavior from various idol fanbases, core fans can formulate an initial strategy for their own group. For example, CF26 and her core fan colleagues examine tutorials from different idol fanbases, selecting and integrating them based on their experience. The core fan then guides general fans in improving the idol’s visibility through continual experimentation, eventually tailoring an algorithmic tutorial that best fits their fanbase. This is because fan groups are heterogeneous. The behavior and characteristics of fans from different idol groups can vary significantly, so core fans must carefully understand and adapt to the style of their corresponding general fan group. Moreover, since these are folk understandings, the algorithmic tutorials created by other fan groups are not necessarily effective. Core fans are often skeptical of tutorials from other groups and prefer to develop their own understanding through repeated experimentation and testing.
Improvements in data performance are often perceptible only to core fans, rather than being quantifiable through objective social media metrics. They observe whether the manipulations lead to higher comment rankings and visibility than other fanbases. Recently, CF26 got information from a fan working at Sina (Weibo company) about how to make a topic become a trending topic. When a topic related to an idol enters the trending section by the recommended algorithm, it gains significant visibility, making this achievement extremely important for fans.
Within a single topic, there are often many posts, but how to intervene so that the algorithm recommends the topic and pushes it into Weibo’s trending list, thereby gaining greater visibility, is a common concern for fans. CF26 mentioned that the new information from the Sina staff asks the fans to engage with the post displayed on the top of the topic, called “source of trend.” This means that interacting with the top posts within a topic—by giving them more retweets, likes, and comments—will make the topic more likely to enter the trending section rather than interacting with other, less prominent posts. When CF26 and her followers tried to make a new trend topic for their idol, they utilized this information to make a new tutorial, and engaged with the “source of trending.” After the topic successfully becomes a new trending topic, CF26 uploads a new post: “The new tutorial is so useful! It’s been trending for about two hours!” CF26 told us that she uses this period as a verification period to ensure the new tutorial makes sense.
4.2.2. Expand Folk Understanding of Algorithms across Platforms in Different Nations.
If such collective action involves multiple countries, people must be able to understand the differences in platform and algorithm policies across these countries. It is important to note that experiences manipulating algorithms can vary across different nations. For instance, in the context of the Korean entertainment industry and social media, Korean fans often engage in a specific strategy of “building within a building,” meaning “commenting within comments,” a layered commenting approach to manipulating data performance.
CF26 highlighted the unique dynamics between Korean and Chinese social media, noting, ”The context in Korean social media differs from that in China, where the regulations are more stringent.” The ”die-hard fan” status, a Weibo feature in China, indicates a high level of interaction between followers and specific bloggers, requiring engagement for at least five days within the past 30 days. This status must be earned separately for each blogger. In China, the ”building within a building” commenting strategy is significantly influenced by the user’s status as a ”die-hard fan” and their VIP identity on Weibo. CF26 explained that if users meet both criteria, their fan group’s comments gain enhanced visibility. However, if they lack the ”die-hard fan” status, the strategy can actually have a detrimental effect.
Additionally, CF26 elaborated that most fans in China do not meet the ”die-hard fan” and VIP criteria. ”die-hard fan” represents a high frequency of interaction with a specific blogger, while achieving VIP status involves monetary contributions. General fans also find it difficult to consistently achieve ”die-hard fan” status across multiple bloggers. As a result, the ”building within a building” strategy may decrease visibility if those who cannot fulfill both requirements conduct algorithmic actions.
In contrast, Korean social media platforms usually do not have the same requirements for enhanced visibility, such as the ”die-hard fan” status and VIP identity. This means that anyone can use the ”building within a building” strategy in Korea to increase visibility without needing a high level of interaction or monetary contributions. The effectiveness of this strategy in Korea, regardless of a user’s status, highlights the different ways platform management policies are implemented in the two countries. While Chinese platforms have stricter rules that limit who can successfully use this strategy, Korean platforms allow all users to benefit from it equally.
CF43 agreed that in China, core fans typically advise against general fans using the ‘building within a building’ comment strategy. CF43 elaborated,
“This is because when general fans, who don’t meet the aforementioned criteria, attempt such commenting, the platform may classify it as an action of a ‘water army,’ perceived as disruptive to social media order. Once labeled as a ‘water army,’ the fan comment is de-prioritized and relegated to the bottom of the comment section, losing all visibility and rendering previous efforts futile.” (CF43)
This difference between the two countries also demonstrates that if core fans do not develop a precise understanding of the algorithms specific to each platform, their collective actions are unlikely to succeed. Without this knowledge, they cannot determine which behaviors will be effective or ineffective and may even risk causing negative consequences.
4.2.3. Expand Folk Understanding of Algorithm across Different Fan Bases.
If this collective action involves different groups of people, it is essential to understand the unique characteristics of these groups and leverage these traits to effectively organize them. Consequently, the interpretation of algorithms and the formulation of ttutorials for algorithm manipulation by core fans vary significantly between cultures and fanbases, and are largely based on empirical observations.
CF11, who manages a data group in an idol’s fan club, highlighted how fans from different countries, even those of the same idol, employ distinct strategies on platforms like YouTube. CF11 noted that view counts of fancams on YouTube are crucial for K-pop idols, and K-pop fans worldwide, regardless of their country, strive to increase these counts. However, CF11 pointed out a stark contrast in techniques: while international fan organizations advise fans to remain logged in while viewing, Chinese fans do not recommend this approach. CF11 further explained their understanding of YouTube’s view count algorithm: “YouTube’s algorithm limits the number of view counts to 3-5 plays per video per day from a single user. Users logged in can’t contribute more than five views. Therefore, to exceed this limit, one should not log in.” CF11 believes these different approaches stem from the varying user habits of the two fan bases,
“Chinese fans are used to repeatedly engaging in data labor, often exceeding five views per day, hence our preference for not logging in. In contrast, the international fan base, predominantly from Europe and the U.S., tends to view 3-5 times daily. For them, ensuring each view is counted becomes vital, making logged-in viewing more appropriate.” (CF11)
4.2.4. Expand Folk Understanding of Algorithm across Different Levels of Algorithm Transparency.
Since this is a collective action targeting algorithms, it must take into account the differences between the algorithms of various platforms, particularly in terms of transparency. Effectively leveraging platform transparency can enhance the efficiency of the action. Even when dealing with platforms with low transparency, people must strive to understand the underlying algorithmic mechanisms. While deciphering algorithms can be intricate and challenging, as seen with platforms like Weibo, some platforms offer more transparency. CF39 mentioned that many platforms provide official explanations of their recommendation systems, simplifying the process for core fans to create tutorials.
For example, BiliBili, a popular platform in China, officially states that the same video can be counted 3-5 times per day per person for play data. Although full playback isn’t necessary for counting views, the completion rate in playing is a significant factor in the recommendation algorithm. This clarity gives fans straightforward tutorials for contributing to BiliBili’s play count.
However, according to our participants, the criteria for BiliBili’s trending list are less clear for fans. Within the Chinese fandom, it’s rumored that a video must achieve 100k plays within 24 hours, along with a specific number of likes, shares, comments, and other metrics, to feature on the trending list algorithmically. Yet, there are instances where videos have made it to the trending list with fewer than 100k plays or exceeded this threshold without being featured. CF39 expressed frustration over this ambiguity, stating,
”We understand that activities like plays, likes, favorites, comments, and shares positively impact trending potential, but the exact mechanism is unclear. We cannot definitively influence the recommendation algorithm. Our only recourse is for fans to engage in data labor blindly and keep experimenting.” (CF39)
This scenario highlights the often opaque and unpredictable nature of social media algorithms, leaving even experienced core fans like CF39 to resort to trial-and-error methods.
We have described how core fans gradually develop a folk understanding of how algorithms operate on different platforms. It is important to note that while these understandings are not entirely conveyed to general fans, they serve as a crucial foundation for core fans to develop algorithm manipulation tutorials. All the core fans we interviewed acknowledged that without a clear understanding of platform algorithms, collective algorithmic action would be impossible. CF34, a core fan with over ten years of experience, explained to us,
”Ten years ago, we lacked understanding of algorithms and the idea of collective algorithmic action. Fans didn’t unite. Even after recognizing the importance of recommendation algorithms, it took time to learn how to coordinate effectively. Without clear direction, mobilizing millions of fans would have been impossible. Now, our actions are organized and disciplined, and fans know precisely what is expected of them, resulting in a significant impact.” (CF34)
In short, having a systematic approach that allows for iterative refinement of communities’ understanding of algorithms and the development of detailed algorithmic theories tailored to the characteristics of specific groups across different platforms and countries is a critical step in organizing collective algorithmic action.
4.3. “Enable the Foolish to Learn, Motivate the Lazy to Act”: The Principles of Mobilizing Millions of Fans (RQ3)
Once core fans are motivated to organize collective fan actions and have gained a sufficiently clear understanding of the underlying mechanisms of algorithms at a folk level, the next challenge is how to effectively mobilize a larger number of general fans to participate in these collective actions. In essence, the folk decoding of algorithms represents merely the initial phase in the algorithmic practices of fans, aimed at enhancing their favorite idol’s data performance. The predominant challenge for most core fans is convincing a wider fan base to participate in this collective action for their idols.
CF37 challenged the previous Chinese media’s portrayal of fans as a highly efficient, organized group that immediately responds to requests for collective algorithmic action. According to CF37, this depiction is misleading. CF37 explained that each act of data engagement—likes, shares, or comments—represents not just a digital footprint but also the genuine affection of a real person behind the screen. As CF37 noted, ”We, all the fans, including core fans like myself, partake in this data labor purely out of love, without any financial or other incentives. Our sole aim is to elevate our idol’s stature.”
However, because this work is unpaid, many fans hesitate to dedicate their time and energy to such tasks. CF37 pointed out that core fans must rely on persuasion, emphasizing both the importance of the data and the simplicity of their tutorials to encourage wider participation in these practices. By framing data engagement as an accessible and meaningful way to support the idol, they hope to motivate more fans to join the collective effort.
According to the 43 core fans we interviewed, to encourage an increasing number of general fans to participate in these algorithmic actions, they have undertaken two primary strategies:
First, they focus on convincing general fans of the significance of fan-driven collective algorithmic actions.
Second, they aim to create straightforward, user-friendly tutorials that are easy to understand and implement for general fans.
4.3.1. Convincing Skeptical General Fans
All core fans in our cases reported they heard skepticism from general fans, who often question the effectiveness of their data labor: ”Is it really useful for me to do this data? Is the time and effort I spent really rewarding my idol accordingly?”
As GF4, a general fan with two years of experience whom we interviewed, put it,
”Participating in algorithmic actions is really exhausting and complicated. I have to figure out what I need to do on each platform, and I spend a lot of time browsing posts and even playing videos, which drains my energy and ties up my devices. Sometimes, I don’t even know what my actions mean—does my small effort as a single user really make a difference? And if I don’t do it, there will be other fans who will.” (GF4)
All the core fans we spoke to believe their primary task, before presenting any tutorial, is to convince other fans of the importance of participating as data labor. This kind of persuasion is key to reducing the free-rider effect among general fans and encouraging active participation in collective algorithmic actions. Core fans inform general fans and make numerous posts about the direct benefits that fan data labor can bring to the idols, often citing examples of other idols’ success stories.
For instance, core fans will argue that if their beloved idol’s data performance is outstanding, similar benefits achieved by predecessors can be expected. If the benefits are not immediately evident from past cases, core fans will highlight the authoritative nature of the platform where they engage in data labor. They also emphasize the prominence of idols who are part of the same data performance competition, thereby crafting a narrative that suggests brands and TV programmers closely monitor this data. This way, core fans imply to the general fans that their idol could gain significant resource boosts through this data and their collective labor efforts.
4.3.2. Making Algorithmic Manipulation Tutorials as Simple as Possible
Moreover, CF16, a core fan who frequently creates algorithmic manipulation tutorials for their fan group, emphasized the importance of making these tutorials as simple as possible for general fans. She explained that this often requires creative approaches to fit seamlessly into fans’ daily routines.
For example, CF16 explained that she developed two tutorials for boosting video plays on Bilibili: one for the mobile app, which requires fans to manually click every 20 minutes, and another for the web version, which can loop the video continuously with no further input, making it much easier and less demanding. Furthermore, CF16 observed that it is more effective to ask fans to boost video plays in the morning when they are at work and can use the web version on their computers. In the evening, fans are more likely to be at home using the mobile app on their smartphones, which is more demanding. As a result, she strategically times her calls for engagement during morning hours to maximize participation.
There are many instances where the tools available for fan activities differ depending on the device or platform used. For example, an auto-tapper app for sending love taps during live idol streams is only available to Android users, while Apple users must manually tap their screens at intervals. As a result, this task is mostly handled by Android users, as it’s more cumbersome for those with Apple devices. Similarly, a Weibo-based tool named ”SHARE”, which simplifies ”controlling comments” with a keyword filtering function, is also limited to Android devices. This led to Android users taking the lead in managing and evaluating content related to their idols in a long period.
However, CF16, CF19, CF36, CF42, and CF43 noted that this changed in 2021 when the SHARE app was reportedly taken down due to infringement issues with Weibo, leveling the playing field between Android and Apple users. This shift allowed Apple users to become more involved in ”controlling comments,” reflecting the evolving landscape of fan algorithmic activities and the tools they use.
In summary, the key to an effective algorithmic tutorial lies in its simplicity and understandability. The complexity of these tutorials is often tailored to the specific fanbase, with their loyalty and dedication playing a crucial role. As CF40, a fan who has been a core fan in various fan groups, explained,
”Core fans for idols with highly devoted fanbases typically employ more complex but detailed tutorials. In contrast, core fans for actors whose fans are mostly casual supporters tend to opt for simpler tutorials, sacrificing some level of detail to attract more participation in data labor.” (CF40)
Although this observation is not backed by quantitative data, there is a consensus among all the core fans we interviewed: The less manual intervention a tutorial requires, the more likely fans will use it for data manipulation. If the general fan can engage in algorithmic action through a straightforward, automated setup on their phones or computers, their willingness to participate is generally high. Conversely, tasks requiring more cognitive engagement and physical effort are met with lower enthusiasm.
P42 concluded in one sentence, “In all, as core fans, we need to enable the foolish to learn while motivating the lazy to act.”
4.4. ”Our Idol Only Has Us!” The Fan Algorithmic Emotionality of Collective Algorithmic Action (RQ3)
Rational persuasion alone is often insufficient to fully mobilize general fans. What supports large-scale mobilization is a powerful set of emotionality, which is widely recognized and continually reinforced within the fan communities. In their efforts to motivate ordinary fans towards collective algorithmic actions, core fans often employ emotional elements alongside objective explanations of the data’s importance. This strategy evokes emotional sympathy, encouraging fans to participate in collective algorithmic actions. Within the fanbase, this type of rhetoric is called ”Guilt-Tripping Fans.”
”Guilt-Tripping Fans” involves conveying to fans that their idol is facing significant challenges and relies solely on their support, thus creating a sense of urgency and need. CF8, a core fan, describes this approach:
”It’s crucial for fans to understand the gravity of the situation, both rationally and emotionally. Our idols need us, not just in a figurative sense. Particularly in China, fans invest financially to fulfill a sense of obligation rather than merely seeking happiness. Therefore, fostering this sense of responsibility is key to compel fans to engage in data performance.” (CF8)
”Guilt-Tripping Fans” remains a contentious topic within the Chinese fandom community. This rhetoric is often seen as exploiting fans’ emotions and coercing them into becoming data laborers. As GF8, a general member of the fanbase, observed, ”In the fan culture, any action that benefits our idol is regarded as a necessity.” She posits that this principle is powerful enough to influence fan behavior in various situations, akin to the effects of emotional coercion.
CF39 elaborated on the strategic use of rhetoric in the context of ”Guilt-Tripping Fans” and fans’ algorithmic emotionality. CF39 said, ”Effective ’Guilt-Tripping’ rhetoric involves storytelling about our idol’s past, present, and future.” CF39 further explained that the past rhetoric focuses on their challenges and perseverance, prompting fans to consider whether they want their idol’s past efforts to be wasted due to their apathy. CF39 added, ”The present rhetoric paints a picture of their current isolation and vulnerability.” CF39 emphasized that it highlights the idol’s lack of support apart from fans, in an environment where company backing and capital often favor those with privileged connections, and external threats are constant. The message is that if even fans abandon them, their situation will become even more dire. CF39 continued, ”The future rhetoric is about the potential outcomes of our actions or inactions.” CF39 elaborated that these scenarios are made as alarming as possible to compel everyone to act, aiming to prevent the idol from falling into unfavorable circumstances.
CF39 concluded that this narrative creates an atmosphere of guilt and regret, prompting fans to imagine the sadness they would feel if their idol were to suffer a tragic fate due to fan negligence. Once fans are convinced by this rhetoric, particularly when it is promoted by influential fans, they are likely to be swayed by the emotions embedded in collective algorithmic actions, effectively turning them into data laborers dedicated to these algorithmic efforts.
However, this ”Guilt-Tripping Fans” rhetoric does not work for all general fans. CF37 points out that since the inflammatory rhetoric used by core fans is not mandatory, they cannot persuade every fan to act in the same way. ”Of course, some general fans are dissatisfied with this, thinking, ’As long as I’m happy, I don’t need to worry about how to help my idol,’” CF37 adds,
”But we actually have to accept the existence of such general fans, as they are only a small fraction of the fan base and do not affect the overall algorithmic emotionality within the fandom. This blending of algorithmic labor with emotionality towards loyalty and love for idols has already become an important part of fan culture and won’t be disrupted by the minority who do not recognize it.” (CF37)
Furthermore, CF37 notes that in fandom. there is a comprehensive organizational strategy in place to ensure that most general fans involved in the fan community can perform algorithmic tasks. CF37 states, ”General fans certainly have the option not to participate in these algorithmic actions or in the algorithm manipulation that we, as core fans, ask they to do. However, when it comes to distributing gifts considering idols from the fan club to general fans, such as priority access to idols’ concert tickets, these algorithmic labor activities are considered as key evaluation criteria.” Through this organizational strategy, the fan community disciplines its members, making fans who engage in algorithmic work the mainstream while those who refuse become the non-mainstream in fandom, unrecognized by the fan community and left on its periphery. In other words, these general fans, who do not participate in the fan-driven collective algorithmic actions, may continue to like their idols but find it challenging to integrate into the corresponding fan community.
”Guilt-Tripping Fans” resonates with previous research on algorithm-driven fan culture (Yin, 2020), which posits that data performance related to algorithms has become a pivotal object of fan sentiment and even a crucial factor in determining fans’ self-identification and in securing recognition within the fan community. Ultimately, through the continuous emotional rhetoric, general fans cultivate an emotional stance within the algorithmic realm, believing that failure to engage in algorithmic labor or other supportive actions for their idols equates to a betrayal of their loyalty to the idol and a denial of their own identity as fans. Driven by the guilt instilled by ”Guilt-Tripping” and their deep affection for their beloved celebrities, countless fans passionately partake in algorithmic labor.
5. Discussion
5.1. From Individual Folk Theory to Collective Understanding and Action towards Algorithms
Our research contributes to HCI studies by demonstrating how a minority can decode algorithms, develop practical tutorials, and disseminate this algorithmic knowledge to a broader base of general users without specialized expertise. This transferring of folk algorithmic understanding empowers users to collectively manipulate algorithms in their day-to-day practices. This process enhances the study of folk theories of algorithms, expanding it from the dimension of understanding to the dimension of action and from the individual level to the collective level.
This study corroborates the findings of previous algorithm research. Folk theories of algorithms (DeVito, 2021; Duffy and Meisner, 2023; Cotter, 2019; DeVito, 2022) provide the theoretical foundation for the core fans to decode algorithms. Core fans continuously gather public information from platforms, professional folk understandings of algorithms from other fan groups, and gossip. As previous scholars have noted, these algorithmic insights from other groups form an external knowledge base that further refines their folk understanding of algorithms (Ticona and TSAPATSARIS, 2023; Zhang et al., 2023). Core fans also observe changes in visibility following algorithm manipulation within their own fan community to refine their understanding of algorithms. The external knowledge base, combined with ongoing internal folk experiments, has led to an increasingly clear understanding of algorithms within the fan community.
This validation process of algorithmic understanding for fan community, directly manifested in the improvement or not of data performance, is interpreted by core fans in several aspects: firstly, whether the tutorial reflects a correct understanding of the algorithm; secondly, whether this tutorial effectively mobilizes the enthusiasm of general fans. In the end, through such a process, algorithm tutorials are validated and become core fans’ reliable cognitive understanding of algorithms, as well as the practical dependence of a large number of general fans. This theoretical process can guide subsequent research on folk algorithm understanding, resistance, and manipulation, guiding scholars to consider several aspects of complete algorithm interaction, such as whether users have formed a folk understanding of the algorithm, whether this understanding can be operationalized, whether this manipulation has produced effective or erroneous results, whether users have relied on these effects to improve their algorithmic practice, and whether reliable validation has been incorporated into their algorithmic knowledge system.
We must point out that, as repeatedly emphasized in the results section, core fans cannot directly understand the platform’s algorithmic mechanisms or access any quantitative data. They must continuously observe changes in idol visibility, which falls entirely within the realm of folk theories of algorithm (Eslami et al., 2019). This perspective is purely from the user’s viewpoint, not from that of a professional algorithm engineer. Despite constant iterations and updates over time and across platforms, the algorithm remains a black box for fans. These iterations merely provide tutorials for algorithmic actions that fans can engage with. In our view, developing such a clear understanding is crucial for transitioning from understanding to action in folk algorithmic theories, especially in mobilizing large-scale collective action.
Moreover, the collective algorithmic understanding and actions of fans remain centralized, determined, and disseminated by a small number of core fans. Zhang et al. (Zhang et al., 2023) argue that fans’ algorithmic imaginaries are co-constructed, and most fans form their algorithmic imaginaries not only through direct engagement with algorithms but also by assimilating second-hand knowledge learned from others. Our empirical data further indicate that fans’ algorithmic imaginations originate from core fans, who play a decisive role in spreading these imaginations to countless general fans and persuading them to engage in algorithmic practices. This finding emphasizes the inequality nature of algorithmic imaginaries within group communication, urging future research to focus more on the genesis and dissemination of algorithmic imaginaries within groups, including interactions between different circles within the group, rather than vaguely understanding it as a part of group culture.
5.2. From Failed or Small-Scale Actions to Successful Large-Scale Computer-Supported Collective Actions
In previous HCI research, the focus has rarely touched on such large-scale, successful computer-supported collective actions. Our study explores a unique case where millions of or even tens of millions of fans mobilize into collective action to manipulate platform algorithms. This action occurs across multiple fan communities, spans major social media platforms, and extends across several countries. Our study validates and extends the effectiveness of the five stages in computer-supported collective action proposed by Shaw et al (Shaw et al., 2014).
First, core fans recognized an urgent and critical need for algorithmic actions to significantly boost their idol’s visibility. They clearly understood the importance of social media data for their idol. They also recognized that this issue was crucial for all general fans, forming the basis for mobilizing millions of general fans. Moreover, all core fans acknowledged the significance of data performance generated by fans. Every core fan we interviewed consistently invested effort in making large-scale algorithmic collective actions a reality, driven by their belief in the reliability of algorithmic labor.
Then, to find effective and appropriate solutions, they had to develop detailed folk algorithmic theories robust enough to guide collective algorithmic actions across different platforms and even in different countries. This step is the most challenging for core fans, as without a solid theoretical foundation to guide their actions, the actions themselves would be impossible. They must continuously conduct grassroots experiments, learn from the experiences of other fan communities, and gather various pieces of information to iteratively refine their folk algorithmic understanding.
When organizing and executing actions, mobilizing large-scale fan participation and ensuring they understand how to act poses a significant challenge for core fans. On one hand, they employ emotional rhetoric to continuously reinforce the importance of participating in collective algorithmic actions among general fans. In a sense, this emotional rhetoric has evolved into a form of emotional coercion, becoming an accepted norm within the fan community. As previous fan studies have revealed, algorithms within fan communities are often emotionalized (Zheng and Xiao, 2023). Our research further explains how algorithmic actions become a shared emotionality among fans. This also fills a gap in the existing theory of computer-supported collective action, highlighting that the emotionality inherent in the collective action itself is a critical component of collective consciousness (Etter and Albu, 2021). On the other hand, core fans must create easy-to-understand, user-friendly, and nearly fully automated algorithm manipulation tutorials to minimize the effort and time required from general fans, thereby maximizing their engagement and contributions. These two aspects together reduce the free-rider effect in fan-driven collective algorithmic actions, allowing general fans to participate in such actions with minimal effort while simultaneously internalizing these actions as intense emotionality, even considering them a key part of their fan identity. These two aspects have also been supported in previous collective actions studies (Olson Jr, 1971; Bonini and Treré, 2024; Treré and Bonini, 2024), where organizers often need to provide strong emotional motivation and clear, actionable guidelines to mobilize large groups effectively. In this study, these points are extended to algorithms: users need to form an emotional connection to the algorithm, such as associating it with a sense of responsibility, and they also need a clear understanding of the actions required to influence the algorithm.
In this whole process, iteration becomes a crucial element. This echoes Shaw et al.’s key factors for the success of computer-supported collective action, namely the need for continuous reflection, evaluation, and improvement of actions (Shaw et al., 2014). Core fans must (1) stay attuned to changes in the platform, as the platform constantly adjusts its algorithms to prevent fans from disrupting its order through algorithmic practices; (2) understand the evolution of computational tools, as these tools are also continuously advancing, requiring core fans to strategically leverage the right tools to support collective action; and (3) comprehend the dynamics of their fan community to fully mobilize their followers. In this context, the core fans’ folk understanding of algorithms and their organization of algorithmic collective actions are not static but are continuously iterated and updated to sustain large-scale computer-supported collective actions.
Our study offers valuable insights for future computer-supported collective actions, particularly those focused on algorithmic interventions. It highlights the necessary conditions for mobilizing the collective action of millions of users. It is worth noting that the case we discussed spans multiple fan communities, targeting algorithmic actions across several countries and platforms, demonstrating the replicability of our proposed theory of collective algorithmic action to some extent. As highlighted in our previous results, these fan communities also learn from one another and draw on each other’s experiences to develop their own collective actions. Therefore, we encourage future HCI scholars to consider the potential generalizability of our theory of computer-supported collective algorithmic action to other groups and actions.
6. Limitations and Future Work
This study has several limitations. First, it focuses on fan communities, which are renowned for their intense enthusiasm and high level of engagement with their idols, making them more likely to participate in collective actions than other user groups. Therefore, how to initiate large-scale collective actions among other user groups remains an open question. Second, the ethnographic research was conducted from 2021 to 2023, when fan collective actions had become routine and highly mature. We mainly focused on successful fan communities, making it difficult to observe moments of failure in collective actions and the factors contributing to such failures. We encourage future research to explore how computer-supported collective actions encounter failure, providing more case studies. Third, our study primarily examines fan collective actions related to algorithms, but fans also engage in collective actions in other areas. For example, previous scholars have examined American fans’ collective support for democratic institutions or the Black Lives Matter movement (Duvall and Heckemeyer, 2018; Jung, 2012). We also encourage future scholars to investigate fan behavior in other forms of collective action.
7. Conclusion
In our study, we examine how millions of fans form effective, self-organized, computer-supported collective actions targeting platform algorithms. We highlight a distinct group of fans motivated by their affection for an idol and the aspiration to enhance the idol’s success, which forms the foundation of their algorithmic emotionality. This unique emotionality framework in algorithmic engagement enables both the dedicated fans, known as core fans and the general fanbase to fully participate in algorithmic work. Core fans play a pivotal role in deciphering algorithms, creating comprehensive manipulation tutorials, and motivating general fans to partake in this algorithm labor zealously. To address the reluctance of most fans to invest substantial time and effort, core fans employ specific strategies: they leverage the emotional rhetoric within fan communities to continuously reinforce the algorithmic emotional stance among fans. Furthermore, core fans work to simplify and automate the process of algorithm manipulation, thus making it easier for general fans to participate in this labor-intensive activity.
Appendix A Higher-level Themes
We provide a summary of the higher-level themes we identified in Table 1. These themes broadly correspond to the section headers in Section 4.
| Third-level themes | Second-level themes | Explanation |
| Recognizing the Necessity of Fan-Driven Collective Algorithmic Action | The Impact of Social Media Action on Idol Visibility | Fans understand that their social media actions—such as liking, sharing, and commenting—directly influence their idol’s exposure and visibility. |
| The Importance of Data Performance in Commercial Collaborations | Fans realize that an idol’s ”data performance” plays a decisive role in brand selection and commercial collaborations. Understanding that enhancing data performance through collective actions can help their idol secure more business opportunities makes these fan-driven actions necessary and meaningful. | |
| Relationship Between Fan-Driven Collective Algorithmic Action and Social Media Algorithms | Through practice, fans have discovered that their strategies (such as ”controlling comments”) can influence social media algorithmic outcomes, thereby increasing their idol’s visibility. This understanding of how algorithms function, coupled with the ongoing adjustment of their strategies, further reinforces the perceived necessity of collective algorithmic action. | |
| Competitive Arenas on Digital Platforms | Fans understand that actions on various digital platforms (e.g., trending topics, popularity rankings) are essential battlegrounds for increasing their idol’s visibility. Their activities are not limited to one platform but span multiple digital domains, extending their actions into different levels and emphasizing the need for various collective actions across platforms. | |
| The Theoretical Foundation for Guiding Collective Algorithmic Actions | Empirical Methods in Decoding Algorithms | Core fans use empirical methods, such as trial and error, to decode social media algorithms. They often gather basic tutorials from other fan groups, adapting them based on their own experience and experimentation to formulate strategies tailored to their fanbase. |
| Adapting Strategies to Different Cultural and Social Media Contexts | Core fans recognize the importance of adapting strategies according to cultural differences and platform-specific policies. For instance, they develop distinct approaches for manipulating data on Korean versus Chinese social media platforms. | |
| Platform-Specific Differences in Algorithm Manipulation | Different social media platforms, like YouTube, Weibo, and BiliBili, have unique algorithms that require tailored manipulation strategies. Core fans develop specialized methods based on the specific criteria and constraints of each platform. | |
| Importance of Algorithmic Knowledge for Effective Collective Action | Core fans emphasize that a clear understanding of platform algorithms is essential for coordinating large-scale, collective algorithmic actions. Without this foundational knowledge, it would be impossible to organize and direct the efforts of millions of fans effectively. | |
| Challenges in Manipulating Opaque and Unpredictable Algorithms | Despite efforts to decode platform algorithms, fans often face challenges due to the opaque and unpredictable nature of social media algorithms, leading them to rely on trial-and-error methods. | |
| Evolution of Collective Algorithmic Actions in Fan Communities | The development of collective algorithmic actions has evolved over time, moving from a lack of understanding to a more organized and disciplined approach, where general fans know what is expected of them, resulting in more impactful actions. | |
| The Principles of Mobilizing Millions of Fans | The Challenge of Mobilizing General Fans for Collective Actions | While core fans are motivated to organize collective actions, the real challenge lies in engaging a larger number of general fans, who may be less committed or lack understanding of the importance of these actions, to actively participate and contribute to the idol’s data performance efforts. |
| Overcoming Skepticism and Reducing the Free-Rider Effect | Core fans face skepticism from general fans who question the effectiveness of their data labor. To overcome this, core fans emphasize the significance of collective algorithmic actions and provide examples of successful outcomes to encourage participation. Core fans also develop easy-to-follow tutorials for general fans to facilitate their engagement in algorithmic actions. These tutorials are designed to be integrated into the daily routines of fans, minimizing the effort required and increasing participation. | |
| Tailoring Tutorials to Device and Platform Accessibility | Core fans create different versions of tutorials based on the specific devices and platforms available to fans, such as mobile apps versus web browsers, to optimize participation across diverse user groups. | |
| Balancing Complexity and Participation | The complexity of algorithmic tutorials is tailored to the dedication level of the fanbase. Tutorials with higher levels of automation tend to have higher participation rates, while more complex tasks requiring manual intervention see lower enthusiasm from fans. | |
| Challenges in Manipulating Opaque and Unpredictable Algorithms | Despite efforts to decode platform algorithms, fans often face challenges due to the opaque and unpredictable nature of social media algorithms, leading them to rely on trial-and-error methods. | |
| The Principle of Effective Fan Mobilization: Simplicity and Motivation | Core fans follow the principle of making learning accessible for all and motivating even the least active fans to participate. The goal is to create tutorials that are simple, automated, and easy to use, thereby maximizing participation in data labor. | |
| The Fan Algorithmic Emotionality of Collective Algorithmic Action | The Limitations of Rational Persuasion in Mobilizing Fans | Rational arguments alone are often insufficient to motivate general fans to participate in collective algorithmic actions. |
| The Strategy of Guilt-Tripping Fans | Core fans use the tactic of ”Guilt-Tripping Fans” to create a sense of urgency and responsibility, thereby encouraging participation in data labor. | |
| Algorithmic Emotionality in Fan Communities | The use of guilt-tripping within the fanbase is seen as a form of algorithmic emotionality, where fans are made to feel obligated to contribute to data labor as a duty to support their idol. |